OpenStudy (anonymous):

Write the equation in standard form for a parabola. Vertex (0,0) , f(0, 1).

OpenStudy (anonymous):

y=x^2+1 or y=-x^2+1

OpenStudy (anonymous):

hmmmmm which one?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I am so sorry that I gave you a wrong answer because there is no equation of parabola can though (0,0) and (0,1) both side. the equation of parabola only has one pint with y-axis, unless, you switch y-axis to x-axis. Then, the equation should be x=±y^2+y. Sorry again.